# Likelihood-Scheduled Score-Based Generative Modeling for Fully 3D PET Image Reconstruction

### v1

@@ VI-D Sample path of likelihood-matched vs fixed updates
Fig. 8 shows the number of likelihood steps taken as a function of generative timestep. Whereas PET-DDS maintains a constant number of likelihood steps per generative timestep (i.e. a horizontal line), our method varies the number of likelihood steps to conform to the likelihood schedule. Relatively fewer steps at the start of the reverse diffusion process wastes less computation, as much information is lost when the re-noising step adds high-variance Gaussian noise to the Tweedie estimate. Fewer steps at the end of the reverse diffusion process reduces the chance of over-convergence to noisy measurement data.
In Section VI-A, our method used a mean of 201 likelihood updates per reconstructed sample (plus 14 for the likelihood scheduling) compared to 2000 fixed likelihood updates for PET-DDS. Whereas PET-DDS can also work well with fewer likelihood updates, it is clear that there are potential efficiency gains to be made by dynamically varying the number of likelihood steps taken.
@@
Figure 8: Number of likelihood steps taken per generative denoising step for a representative 2D PET reconstruction with likelihood-scheduled SGM-based PET reconstruction. For comparison, PET-DDS takes a fixed number of likelihood steps per generative denoising step (reported by Singh et al. from 4 to 20).)
VI-E 2D reconstruction uncertainty
For a fixed Poisson noise realization, we sample reconstructions based on the score-based prior. Fig. 9 shows varied reconstructions from simulated data at two count levels, demonstrating the model’s ability to generate diverse samples. Higher counts result in less shape variation, closer resemblance to the ground truth, and fewer artifacts. The impact of dose level and generative timesteps on SGM performance will be explored in future work.
@@
VI-F Comparison to MLEM for varying iteration
@@
VI-G Real 3D data
VI-G1 Quantitative results
@@
Figure 11: Coefficient of variation (CoV) against PLL for reconstructions from 10% counts of real DPA714 3D measured data. CoV is measured as the mean pixel-wise coefficient of variation across reconstructions from 5 realizations of noisy measured data, as measured in a region of white matter.
VI-G2 Qualitative results
In Fig. 12 we show reconstructions that achieve equivalent likelihood (to 21 iterations of MLEM) with different methods from real 10% count data. We note that our introduction of perpendicular SGMs resolves the slice inconsistency (seen as alternating intensity transverse slices) seen for PET-DDS’s coronal and sagittal slices. Both of the SGM-based methods display lower noise than the MLEM reconstruction, consistent with Fig. 11.
Figure 12: Example reconstructions for real IRIS data in 3D. All columns except “Clinical” used the ParallelProj projector [38] without positron range modeling; “Clinical” reconstructions used Siemens’ proprietary tools. Reconstructions from 10% count data match the PLL corresponding to 21 iterations of MLEM.
VII Discussion
Our results demonstrate that our method has a significantly lower hyperparameter tuning burden than PET-DDS, with the option of just tuning a single hyperparameter $N_{\text{MLEM}}$ to directly vary the balance between the prior and the likelihood. This replaces tuning strength of MAP regularization $\lambda_{\text{MAP}}$; number of MAP iterations per generative step $N_{\text{MAP}}$; strength of RDP regularization $\lambda_{\text{RDP}}$ and gradient ascent step size $\delta$.
Hallucinations in reconstructions are a known issue with SGM-based reconstruction. While hallucinations are present in some of the example low-dose reconstructions shown, it should be noted that this is an expected result; count levels were set deliberately low to explore the setting where structures are not clearly discernible from the OSEM reconstructions (and therefore cannot be easily reconstructed by MAP-EM even when relying on edge preservation priors).
@@ VII Discussion
Further work on 3D modeling with pre-trained SGMs could improve the image quality in 3D further, for example by investigating the use of 2.5D training approaches or latent diffusion models. Lastly, the methods presented may find use cases in other medical image modalities where model-based iterative steps are combined with pre-trained SGMs, such as magnetic resonance imaging or computed tomography [9].
VIII Conclusion

### v2

@@ 6.4 Sample path of likelihood-matched vs fixed updates
Fig. 8 shows the number of likelihood steps taken as a function of generative timestep. Whereas PET-DDS maintains a constant number of likelihood steps per generative timestep (reported by Singh et al. from 4 to 20 [12])), our method varies the number of likelihood steps to conform to the likelihood schedule. Relatively fewer steps at the start of the reverse diffusion process wastes less computation, as much information is lost when the re-noising step adds high-variance Gaussian noise to the Tweedie estimate. Fewer steps at the end of the reverse diffusion process reduces the chance of over-convergence to noisy measurement data.
In Section 6.1, our method used a mean of 201 likelihood updates per reconstructed sample (plus 14 for the likelihood scheduling) compared to 400-2000 fixed likelihood updates for PET-DDS (depending on the number of likelihood steps per generative timestep, reported from 4 to 20 [12]). It is clear that there are potential efficiency gains to be made by dynamically varying the number of likelihood steps taken.
@@
Figure 8: Number of likelihood steps taken per generative denoising step for a representative 2D PET reconstruction with likelihood-scheduled SGM-based PET reconstruction.
@@
6.5 Effect on lesion recovery
A major strength of unsupervised diffusion model approaches relative to supervised approaches such as FBSEM-net is their ability to resolve out-of-distribution features such as lesions [12]. To test this, we inserted a hot lesion into a 2D test dataset on the boundary between gray and white matter. We then performed reconstruction with this out-of-distribution dataset to investigate the lesion recovery performance of our approach. Figures 10 and 11 verify that our approach does not adversely affect lesion recovery and that our method matches or outperforms the improvements to the contrast recovery coefficient (CRC) that are claimed by PET-DDS. FBSEM-net-adv (FBSEM-net with an advanced neural network) performs notably poorly at this task, highlighting that unsupervised approaches display greater flexibility to reconstructing out-of-distribution datasets (and thereby achieve superior lesion recovery) compared to supervised approaches such as FBSEM-net.
Figure 10: Evaluation of methods for a lesion recovery task, with an out-of-distribution hot lesion inserted into a test dataset. NRMSE is measured globally, while CRC is evaluated on the lesion itself, with a large area of white matter non-overlapping the lesion chosen as the background region. OSEM and MAP-EM ($\beta=3$) have iteration number varied from 5 to 40 and 10 to 100 respectively. For PET-DDS, $\lambda$ was varied from 0.5 to 2.25, whereas for our method $N_{MLEM}$ was varied from 11 to 18. Results were averaged over 7 Poisson noise realizations. See Fig. 11 for corresponding images.
Figure 11: Visualization of each method’s performance for resolving an out-of-distribution hot lesion inserted into a test dataset. Where hyperparameters were varied, the image for each method was chosen as the image with the best NRMSE for a CRC value of at least 0.55. See Fig. 10 for each method’s corresponding quantitative performance.
6.6 2D reconstruction uncertainty
For a fixed Poisson noise realization, we sample reconstructions based on the score-based prior. Fig. 9 shows varied reconstructions from simulated data at two low count levels. As counts increase, less shape variation, closer resemblance to the ground truth, and fewer artifacts are observed. At both count levels, the mean image appears smoother and better matches the ground truth than the samples. The impact of dose level and generative timesteps on SGM performance will be explored in future work.
6.7 Comparison to MLEM for varying iteration
@@
Figure 13: Example reconstructions for real $[^{18}$]DPA-714 data in 3D. All columns except “Clinical” used the ParallelProj projector [33] without PSF; “Clinical” reconstructions used Siemens’ proprietary tools. Reconstructions from 10% count data match the PLL corresponding to 21 iterations of MLEM.
Figure 14: Coefficient of variation (CoV) against PLL for reconstructions from 10% counts of real $[^{18}$]DPA-714 3D measured data. CoV is measured as the mean pixel-wise coefficient of variation across reconstructions from 5 realizations of noisy measured data, as measured in a region of white matter (selected as a large visually uniform rectangle from the ground truth).
6.8 Real 3D data
6.8.1 Qualitative results
In Fig. 13 we show reconstructions that achieve equivalent likelihood (to 21 iterations of MLEM) with different methods from real 10% count data. We note that our introduction of perpendicular SGMs resolves the slice inconsistency (seen as alternating intensity transverse slices) seen for PET-DDS’s (with a single SGM) coronal and sagittal slices. Both of the SGM-based methods display lower noise than the MLEM reconstruction at this likelihood level. (It should be noted that PET-DDS can eliminate slice inconsistency at lower likelihood at the cost of a blurring effect from the axial RDP.)
6.8.2 Quantitative results
@@
Table 2: Real computing time in seconds for computing a single sample with selected methods in 3D. PET-DDS was evaluated with 4 gradient-based steps per diffusion iteration, while our approach was used with $N_{\text{MLEM}}=21$ (corresponding to images in Fig. 13). Timings include constructing the likelihood schedule with 21 steps of MLEM for our approach.
Method Time (s)
25 SGM steps 100 SGM steps
MLEM 33 33
PET-DDS with single SGM 116 398
PET-DDS with perp. SGMs 137 488
Ours ($\delta=1.0$) 145 385
Ours ($\delta=2.0$) 126 369
Ours ($\delta=4.0$) 116 368
6.8.3 Timing
We report the real computing time for 3D reconstruction with our proposed method and PET-DDS (both with and without the use of perpendicular SGMs) in Table 2. Our approach is slightly more time efficient than PET-DDS for low-dose PET reconstruction (on a fair comparison with the same perpendicular pre-trained SGMs), with the overhead introduced from SGM steps greater than in 2D. The times for our approach include the time taken to construct the likelihood schedule; as this only occurs once per dataset, computing more samples is even more efficient with our method.
7 Discussion
Our results demonstrate that our method has a significantly lower hyperparameter tuning burden than PET-DDS, with the option of just tuning a single hyperparameter $N_{\text{MLEM}}$ to directly vary the balance between the prior and the likelihood. This replaces tuning strength of MAP regularization $\lambda_{\text{MAP}}$; number of MAP iterations per generative step $N_{\text{MAP}}$; strength of RDP regularization $\lambda_{\text{RDP}}$ and gradient ascent step size $\delta$. In particular, this work integrates SGM priors with the standard clinical heuristic and vendor recommendation for the number of MLEM iterations.
Hallucinations in reconstructions are a known concern with SGM-based reconstruction. While hallucinations are present in some of the example low-dose reconstructions shown, it should be noted that this is an expected result; count levels were set deliberately low to explore the setting where structures are not clearly discernible from the OSEM reconstructions (and therefore cannot be easily reconstructed by MAP-EM even when relying on edge preservation priors). While individual samples may vary, given enough samples we can obtain a lower variance mean estimate. Our likelihood scheduling approach also makes it easier to increase the level of consistency with a model-based reconstruction by increasing $N_{\text{MLEM}}$ (and potentially spot hallucinations); it could also be integrated with recent approaches for reducing hallucinations on out-of-distribution data [37].
@@ 7 Discussion
Further work on 3D modeling with pre-trained SGMs could improve the image quality in 3D further, for example by integrating our likelihood scheduling approach with 2.5D training approaches, latent diffusion models or patch-based approaches [38]. Lastly, the methods presented may find use cases in other medical image modalities where model-based iterative steps are combined with pre-trained SGMs, such as magnetic resonance imaging or computed tomography [9].
8 Conclusion